# A Comprehensive Survey on Machine Learning Driven Material Defect Detection: Challenges, Solutions, and Future Prospects

### v1

@@ 2.3. Material Defect Detection
Material defect detection, as an important branch of defect detection, is a crucial area of research in industrial production to prevent safety issues caused by material defects. Rapid and non-destructive detection of potential defects in the material production process is a critical aspect of ensuring material quality. Common material defects can be categorized into surface defects and structural defects. Surface defects typically manifest as localized irregularities on the material surface, including irregular or abnormal features such as cracks, scratches, dents, colour variations, bubbles, wear, foreign objects, and more (zhao2020hole, ; raizada2021surface, ). Structural defects, on the other hand, involve abnormalities in the overall structure and shape of the material, such as shape defects (deformation, twisting, or bending), structural defects (improper connections, joints, or fixations), structural inconsistencies, and defects in connecting parts. Structural defects often exhibit a more complex manifestation in real-world applications and may include surface defects (zhao2023structural, ; li2020strong, ; zhang2020defect, ). Fig. 5, Fig. 5, Fig. 5, and Fig. 5 illustrate some common defect samples found in materials, where the Image column contains the original defect images and the GT column includes the corresponding ground truth labels.
Common methods for material defect detection generally include magnetic particle testing (MPT), penetrant testing (PT), eddy current testing (ECT), ultrasonic testing, X-ray inspection, and machine vision inspection. The relevant introduction and comparison among these detection methods are as below:
  • MPT is a non-destructive testing method used to detect surface and near-surface defects in materials. It relies on the properties of magnetic powders, which are applied to the material surface after the application of a magnetic field. If defects are present, the magnetic particles gather at the defect locations, forming visible lines or spots, thereby indicating the position and shape of the defect. MPT is sensitive and cost-effective, but it is limited to detecting surface defects in ferromagnetic materials and may not be suitable for real-time automated detection (burke2013review, ; lugin2023automated, ).
  • PT is a non-destructive testing method for discovering surface or near-surface cracks, pores, or other open defects. It relies on the penetration behaviour of a liquid into the defect surfaces, revealing hidden flaws in the tested object. PT is applicable to various materials, offering advantages in surface defect detection, but it may have limitations in detecting defects in porous materials, and the detection speed is relatively slow with sensitivity influenced by various factors (kumar2015review, ; mehta2015surface, ).
  • ECT is a non-destructive testing method for detecting surface and near-surface defects in conductive materials. It is based on Faraday’s electromagnetic induction principle, generating eddy currents by introducing an alternating magnetic field to identify defects through variations in eddy currents. ECT is fast, sensitive, and suitable for high-temperature defect detection. However, it may lack accuracy in displaying defect shapes and sizes, and the detection progress can be slow (hughes2014near, ).
  • X-ray Inspection is a non-destructive testing method utilizing X-rays to penetrate materials and measure the material’s absorption of X-rays, providing information about internal structures and defects. This method is versatile, and capable of detecting both surface and internal defects with high efficiency. However, it poses radiation risks to operators, and the equipment is generally complex and expensive, with detection accuracy influenced by environmental conditions (du2019approaches, ).
  • Ultrasonic Testing is a non-destructive testing method using ultrasonic waves to detect internal or surface defects in materials. It is based on the propagation characteristics of ultrasonic waves, sending pulses and receiving reflected signals to identify abnormalities in the material. Ultrasonic testing is simple, efficient, and applicable to both surface and internal defect detection. However, it can be more challenging for complex architectural materials, and its effectiveness is limited for non-conductive materials (honarvar2020review, ).
  • Machine Vision Inspection is a technique utilizing computer vision systems for defect detection and quality control. It involves capturing images and using computer algorithms to analyze and process images to detect, recognize, and evaluate defects in the tested object. Machine vision is versatile, highly accurate, and not significantly influenced by material and defect types, making it suitable for various applications (ren2022state, ).
Figure 2. Textile fabric defects in (silvestre2019public, )
Figure 3. Magnetic tile surface defects in (huang2020surface, )
Figure 4. Commutator surface defects in (Tabernik2019JIM, )
Figure 5. Textured surface defects in (wieler2007weakly, )
@@
Through the above comparison of existing methods for material defect detection, it is evident that machine vision-based defect detection methods have significant advantages compared to other approaches. They enable automated non-destructive testing for virtually any material defect, offering high efficiency, safety, and precision. Importantly, defect detection through machine vision seamlessly integrates with the rapidly evolving field of artificial intelligence, making existing material defect inspection more intelligent. This approach has garnered significant attention and preference from both academia and industry. Therefore, machine vision-based defect detection methods will serve as the central focus of this paper.
@@
The image analysis phase constitutes the core of the entire machine vision defect detection system, as specific defect detection algorithms will be deployed and executed here. For defect detection using traditional unsupervised machine learning algorithms, manually designed defect feature extraction and classification algorithms can be directly executed. If deep learning algorithms are employed for defect detection, steps such as defect image annotation, neural network model design, and model training are necessary, ultimately resulting in a defect detection model that can be deployed and executed. Lastly, the visual detection system records defect feature information outputted by the defect detection model, such as defect category, defect size, defect location, etc., in a database. Necessary actions are then taken for the defective material, such as alarms or removal.
Figure 6. Illustration of material defect detection system based on machine vision.
@@ 2.4. Challenges in Material Defect Detection
Compared to common defect detection tasks, material defect detection technology necessitates consideration of specific industry attributes, material characteristics, and the material production environment. Consequently, this introduces new challenges and complexities to the realm of material defect detection. The current challenges facing machine learning-based material defect detection include:
@@
  • Data Quality Issues: Data quality is crucial for the performance of machine learning models. Low-quality data may introduce noise, leading to misjudgments or missed detections. Improving data quality is a significant challenge in material defect detection using machine learning, as the task requires a substantial amount of annotated data for training. However, practical applications often encounter issues such as data fuzziness and noise interference, impacting model training effectiveness and detection performance.
  • Identification of Complex Defects: Recognizing complex defects poses a significant challenge in machine learning for material defect detection. Efficient feature extraction and classification algorithms are required to accurately identify these defects, extracting useful features to enhance detection performance. The diverse types, shapes, sizes, and positions of defects in materials make identification challenging. Traditional detection methods often struggle to accurately identify complex defects such as cracks and pores.
  • Multimodal Data Processing: Handling multimodal data is a crucial challenge in machine learning for material defect detection. Different data types exhibit distinct characteristics and noise levels, necessitating the development of specialized techniques to address these differences. Additionally, integrating data from various sources to improve detection performance presents a challenge. Material defect detection involves multiple data types, including images, ultrasonic waves, X-rays, etc. Effectively processing these multimodal data, extracting useful features, and enhancing detection performance pose challenges.
  • Real-Time Online Detection: Real-time online detection is a crucial challenge in machine learning for material defect detection. Developing efficient algorithms and optimizing models to reduce computation time and resource consumption are necessary to achieve real-time online detection. Maintaining model performance and accuracy in real-time environments is also challenging. Fast, reliable, and real-time defect detection is essential on production lines or in real-time applications, but some current machine learning methods may struggle to meet these requirements.
  • Interpretability and Transparency: Interpretability and transparency are important challenges in machine learning for material defect detection. Developing new technologies and methods to analyze the internal workings of models is essential to enhance model transparency and interpretability. Striking a balance between high performance and improved model transparency and interpretability is a challenge. For critical application areas like medical devices and safety-related fields, increasing model transparency and interpretability is a crucial requirement. However, current deep learning models are often perceived as ”black boxes,” making it difficult to explain the reasons behind their decisions.
  • Privacy and Security: Privacy and security present important challenges in machine learning for material defect detection. Efficient data encryption techniques and secure storage solutions need to be developed to protect privacy and security. Additionally, ensuring models are not subject to malicious attacks or tampering is essential. Privacy and security issues must be considered when dealing with material defect detection. Challenges include secure data storage and transmission, as well as maintaining the confidentiality of models.
@@ 2.5. Survey Overview
Facing numerous challenges in material defect detection, the rapid development of visual hardware and computer vision technologies, particularly those based on deep learning, has led machine learning-guided defect detection algorithms to gradually dominate research. Researchers have innovatively applied various machine learning algorithms to defect detection, significantly improving accuracy, efficiency, and applicability. This study, centred around defect detection technology, explores the application scenarios and advantages and disadvantages of different material defect technologies. It not only provides a comprehensive literature review and future research directions for subsequent researchers but also offers extensive guidance for industrial practitioners.
This paper summarizes recent hot research technologies applied to material defect detection in Fig. 7 and provides detailed discussion and analysis in the following content. Unsupervised detection algorithms are further categorized into traditional unsupervised machine learning algorithms and unsupervised deep learning algorithms. For situations where defects are relatively simple and defect types are known, handcrafted defect features can be used, and rapid defect detection can be achieved using traditional unsupervised machine learning algorithms. For defects without manual annotations, unsupervised deep learning algorithms are needed to automatically learn defect features from defect samples and accurately differentiate between normal and defective samples. Additionally, in cases where a sufficient number of annotated defect samples are available, supervised learning algorithms can be employed. Generally, deep neural network models are used for self-learning features, enabling defect detection, classification, and segmentation. However, in realistic industrial scenarios, annotated defect samples are often scarce. Thus, the latest semi-supervised or weakly supervised learning algorithms are applied to make reasonable learning decisions for labelled and unlabeled defect samples. In situations where defect samples are extremely scarce or the ratio of defect to normal samples is highly imbalanced, modern generative learning methods can be applied for the artificial augmentation of defect samples, balancing the disparity in the quantity of positive and negative samples. For learning new defect samples, reinforcement techniques can also be applied. Furthermore, to meet the industrial detection criteria for real-time and high detection efficiency, model compression techniques are actively applied during the deployment stage to enhance the operational efficiency of defect detection models. Although machine learning-based material defect detection has achieved overall improvements in detection accuracy and efficiency, challenges such as the complexity of defect types, generalization issues in defect detection models, and high costs in model training and deployment still exist.
@@ 3.1. Unsupervised Learning Techniques in Material Defect Detection
This section elaborates on the unsupervised learning techniques employed in material defect detection. Fig. 8 outlines its general workflow comprising four main steps: data collection, data preprocessing, feature extraction, and anomaly detection. Initially, a sensor captures data, forming the original dataset. This dataset is then preprocessed by removing outliers, normalizing values, and filling in missing data to produce a processed dataset. Key features are extracted using methods such as PCA, histograms, and DBNs. Finally, anomaly detection is performed using techniques like clustering, SVM, and classification networks to identify and classify material defects.
Figure 8. Illustration of the general working process for conducting the material defect detection with unsupervised learning techniques.
3.1.1. Histogram-based Techniques
Histogram-based techniques have widely been used in material defect detection (kumar2021histogram, ). The method involves dividing a wood surface image into equal-sized blocks, analyzing each block’s colour using a percentile colour histogram, and extracting texture features through singular value decomposition (SVD) (song2015wood, ). These features are combined into a feature vector for each block. A support vector machine (SVM) classifier is then used to identify defects, particularly effectively detecting knot-type defects on wood surfaces. An improved Otsu method, Weighted Object Variance (WOV) (yuan2015improved, ), is proposed for efficient surface defect detection in machine vision. WOV adjusts the thresholding technique, traditionally used in image processing, to better handle varying defect sizes and histogram types. It ensures more accurate segmentation, leading to higher defect detection rates and lower false alarms. Experimental results demonstrate WOV’s superiority over traditional methods like Otsu and maximum entropy in detecting defects on various surfaces. Erazo-Aux et al. (erazo2019histograms, ) introduce a new method for detecting defects in thermal images of composite materials using pulsed thermography. It uses local histograms of oriented gradients for simple yet effective differentiation of defects from the material background without relying on image enhancement or complex analysis. The method demonstrates high accuracy, particularly in images of carbon fibre-reinforced plastic, achieving robust detection with an average AUC score of 0.949. This approach is more straightforward and less costly than other methods requiring extensive preprocessing. In order to detect the defects in sewers, Mahmoud et al. (halfawy2014automated, ) present a pattern recognition algorithm designed to automate defect detection and classification in sewer pipes using CCTV inspection images. It employs histograms of oriented gradients (HOG) and a support vector machine (SVM). The process involves image segmentation to identify potential defect areas, followed by SVM classification using trained HOG features.
3.1.2. Clustering Algorithm
Clustering algorithms are a class of unsupervised machine-learning techniques that can effectively detect material defects (jin2020machine, ). These algorithms group similar data points together based on certain features, and this characteristic makes them suitable for identifying anomalies or defects in materials. Thomas et al. (leung2001representing, ) focus on and develop a method to identify materials like concrete, rug, marble, or leather based on their textures, focusing on surface reflectance and normals. We introduce ”3D textons,” small prototype surface patches with unique geometric and photometric properties, such as ridges or grooves. Each texton is associated with an appearance vector, capturing light distribution under various conditions. Using a clustering technique, we create a compact library of these textons. This allows for the characterization and recognition of materials from a few images, even under novel lighting and viewing conditions. The method also predicts how materials appear in different environments. In high-tech materials such as semiconductor manufacturing, a spatial defect diagnosis system for semiconductor manufacturing has been proposed to estimate defect clusters and their separation, including both convex and non-convex types. The approach (wang2008recognition, ) uses a spatial filter to clean wafer maps and a hybrid of entropy fuzzy c-means and spectral clustering to isolate defect patterns. A decision tree based on cluster features helps identify defect types. Tested with data from a DRAM company in Taiwan, the method effectively extracts and classifies various defect patterns, showing potential for application in other industries like liquid crystal and plasma display manufacturing. For steel materials, unsupervised machine learning was used to classify images from the Northeastern University Steel Surface Defects Database, which contains micrographs of different defects in hot-rolled steel (cohn2021unsupervised, ). The study employs the VGG16 convolutional neural network, pre-trained on the ImageNet dataset, to extract features from the micrographs. The principal component analysis is then applied to refine these features, followed by k-means clustering for image classification without requiring labelled data. The method achieves an impressive 99.4% accuracy and can classify new images without retraining. This represents a significant improvement over previous studies. Additionally, a sensitivity analysis is conducted to understand the impact of each step on the classification performance, offering valuable insights for applying unsupervised machine learning in materials science.
Unlike K-means clustering, Hierarchical Clustering builds a hierarchy of clusters through a bottom-up approach (agglomerative) or top-down approach (divisive). Jorge et al. (igual2020hierarchical, ) presents a hierarchical clustering approach using impact-echo techniques for defect detection in materials. Feature vectors, indicating wave propagation changes caused by defects, are analyzed as mixtures of Gaussians. Clusters are merged using the Kullback-Leibler divergence. Applied to aluminium alloy specimens, the method successfully classifies them into categories like no defect, one hole, one crack, and multiple defects, effectively distinguishing between defective and non-defective materials. In addition, Brzakovic et al. (brzakovic1990approach, ) introduce an expert system for identifying and categorizing defects in materials with complex textures using digitized images. It involves a two-step image segmentation process and a material-specific defect classification approach. A hierarchical classification scheme incorporating shape, size, and texture descriptors is used for wood. The system, tested on parquet samples, has successfully detected and classified various defects.
3.1.3. Principal Component Analysis

### v3

@@ 2.3. Material Defect Detection
Material defect detection (MDD), as an important branch of defect detection, is a crucial area of research in industrial production to prevent safety issues caused by material defects. Rapid and non-destructive detection of potential defects in the material production process is a critical aspect of ensuring material quality. Common material defects can be categorized into surface defects and structural defects. Surface defects typically manifest as localized irregularities on the material surface, including irregular or abnormal features such as cracks, scratches, dents, colour variations, bubbles, wear, foreign objects, and more (zhao2020hole, ; raizada2021surface, ). Structural defects, on the other hand, involve abnormalities in the overall structure and shape of the material, such as shape defects (deformation, twisting, or bending), structural defects (improper connections, joints, or fixations), structural inconsistencies, and defects in connecting parts. Structural defects often exhibit a more complex manifestation in real-world applications and may include surface defects (zhao2023structural, ; li2020strong, ; zhang2020defect, ).
@@
Common methods for material defect detection generally include magnetic particle testing (MPT), penetrant testing (PT), eddy current testing (ECT), ultrasonic testing, X-ray inspection, and machine vision inspection. Machine vision-based defect detection methods have significant advantages compared to other approaches. They enable automated non-destructive testing for virtually any material defect, offering high efficiency, safety, and precision. Importantly, defect detection through machine vision seamlessly integrates with the rapidly evolving field of AI, making existing material defect inspection more intelligent. This approach has garnered significant attention and preference from both academia and industry. Therefore, machine vision-based defect detection methods will serve as the central focus of this paper.
@@
The image analysis phase constitutes the core of the entire machine vision defect detection system, as specific defect detection algorithms will be deployed and executed here. For defect detection using traditional unsupervised ML algorithms, manually designed defect feature extraction and classification algorithms can be directly executed. If DL algorithms are employed for defect detection, steps such as defect image annotation, neural network model design, and model training are necessary, ultimately resulting in a defect detection model that can be deployed and executed. Lastly, the visual detection system records defect feature information outputted by the defect detection model, such as defect category, defect size, defect location, etc., in a database. Necessary actions are then taken for the defective material, such as alarms or removal.
Figure 2. Illustration of a material defect detection system based on machine vision.
2.4. Experimental Evaluation and Benchmarks
Benchmarking methods for MDD requires comprehensive experimental evaluations using standardized datasets and metrics to ensure reproducibility and enable fair comparisons across diverse defect types and complexities. Widely used datasets, such as NEU (song2014automatic, ), DAGM 2007 (wieler2007weakly, ), and Severstal (severstal2019, ), along with others listed in Table 3, provide annotated defect images spanning various material categories, including steel and metal, fabric, texture, concrete, and industrial materials. These datasets facilitate targeted assessments, with NEU (song2014automatic, ), RSDDs (Gan2017, ), and UCI Steel Plates (UCI_SteelPlatesFaults, ) focusing on steel defects, AITEX (SilvestreBlanes2019, ) and Tianchi Cloth (Qi2023, ) addressing fabric issues, and texture-related datasets like MVTec AD (bergmann2019mvtec, ) and TILDA (TILDA, ) covering surface patterns. Metrics such as accuracy, precision, recall, F1-score, and inference time are crucial for evaluating model performance, providing insights into overall effectiveness, robustness against errors, and suitability for real-time applications. By leveraging these datasets and metrics, researchers can perform standardized evaluations, fostering innovation, enabling fair comparisons, and identifying areas for improvement in defect detection methods.
Table 3. Overview of common datasets for material defect detection
Material Category Dataset Link
Steel and Metal NEU (song2014automatic, ) http://faculty.neu.edu.cn/yunhyan/NEU_surface_defect_
database.html
RSDDs (Gan2017, ) http://icn.bjtu.edu.cn/Visint/resources/RSDDs.aspx
UCI Steel Plates (UCI_SteelPlatesFaults, ) https://archive.ics.uci.edu/ml/datasets/Steel+Plates+Faults
Severalst (severstal2019, ) https://www.kaggle.com/c/severalst-steel-defect-detection/data
GC10-Det (Lv2020, ) https://www.kaggle.com/datasets/alex000kim/gc10det
Fabric AITEX (SilvestreBlanes2019, ) https://www.aitex.es/afid/
Fabric Stain (FabricStainDataset2023, ) https://www.kaggle.com/priemshpathirana/fabric-stain-dataset
Tianchi Cloth (Qi2023, ) https://tianchi.aliyun.com/competition/entrance/
231666/information
Texture and Miscellaneous TILDA (TILDA, ) https://hci.iwr.uni-heidelberg.de/node/3616
Kylberg (Kylberg2011, ) http://www.cb.uu.se/ gustaf/texture/
DAGM 2007 (wieler2007weakly, ) https://hci.iwr.uni-heidelberg.de/node/3616
MVTec AD (bergmann2019mvtec, ) https://www.mvtec.com/company/research/datasets/mvtec-ad
Concrete and Building SDNET 2018 (Dorafshan2018, ) https://www.kaggle.com/aniruddhsharma/structural-defects
-network-concrete-crack-images
Surface Crack (Ozgenel2018, ) https://www.kaggle.com/arunrk7/surface-crack-detection
Bridge Cracks (Li2019, ) https://github.com/lsyksjr/Bridge_Crack_Image_Data
Industrial Kolektor (Tabernik2020, ) https://www.vicos.si/Downloads/KolektorSDD
DeepPCB (Tang2019, ) https://github.com/tangsanli5201/DeepPCB
elpv-dataset (Buerhop-Lutz2018, ) https://github.com/zae-bayern/elpv-dataset
NanoTWICE (Carrera2017, ) https://www.mi.imati.cnr.it/ettore/NanoTWICE/
@@ 2.5. Challenges in Material Defect Detection
Compared to common defect detection tasks, MDD technology necessitates consideration of specific industry attributes, material characteristics, and the material production environment. Consequently, this introduces new challenges and complexities to the realm of MDD. The current primary challenges facing ML-based MDD include:
@@
  • Identification of Complex Defects: Recognizing complex defects poses a significant challenge in ML for MDD. Efficient feature extraction and classification algorithms are required to accurately identify these defects, extracting useful features to enhance detection performance. The diverse types, shapes, sizes, and positions of defects in materials make identification challenging. Traditional detection methods often struggle to accurately identify complex defects such as cracks and pores.
  • Multimodal Data Processing: Handling multimodal data is a crucial challenge in ML for MDD. Different data types exhibit distinct characteristics and noise levels, necessitating the development of specialized techniques to address these differences. Additionally, integrating data from various sources to improve detection performance presents a challenge. MDD involves multiple data types, including images, ultrasonic waves, X-rays, etc. Effectively processing these multimodal data, extracting useful features, and enhancing detection performance pose challenges.
  • Real-Time Online Detection: Real-time online detection is a crucial challenge in ML for MDD. Developing efficient algorithms and optimizing models to reduce computation time and resource consumption are necessary to achieve real-time online detection. Maintaining model performance and accuracy in real-time environments is also challenging. Fast, reliable, and real-time defect detection is essential on production lines or in real-time applications, but some current ML methods may struggle to meet these requirements.
@@ 2.6. Survey Overview
Facing numerous challenges in MDD, the rapid development of visual hardware and technologies, particularly those based on DL, has led ML-guided defect detection algorithms to gradually dominate research. Researchers have applied various ML algorithms to defect detection, significantly improving accuracy, efficiency, and applicability. This study, centred around defect detection technology, explores the application scenarios and advantages and disadvantages of different MDD technologies. It not only provides a comprehensive literature review and future research directions for subsequent researchers but also offers extensive guidance for industrial practitioners.
This paper summarizes recent research advancements in MDD, as illustrated in Fig. 3, and provides detailed discussion and analysis in the subsequent sections. Our classification framework emphasizes the practical relevance and unique contributions of each technique in defect detection, categorizing them into unsupervised learning (UL), supervised learning (SL), semi-supervised learning (SSL), reinforcement learning (RL), and generative learning (GL). This distinction reflects the specialized roles these methods play in addressing diverse defect detection challenges.
UL algorithms are divided into traditional ML and DL methods. For scenarios where defect types are relatively simple and pre-defined, traditional unsupervised ML algorithms can utilize handcrafted features for rapid defect detection. For more complex cases without manual annotations, unsupervised DL methods are required to automatically learn features and distinguish defective samples from normal ones. GL, although often classified as a subset of UL, is treated separately in our framework due to its transformative role in addressing data scarcity. Techniques such as GANs and VAEs are instrumental in generating synthetic defect samples, balancing the sample distribution, and improving model robustness.
In situations where sufficient annotated samples are available, SL algorithms, particularly deep neural networks, can perform feature self-learning for defect detection, classification, and segmentation. However, given the scarcity of annotated samples in real-world industrial scenarios, semi-supervised or weakly SL methods are increasingly applied, enabling effective utilization of both labeled and unlabeled samples. RL is highlighted as a standalone category due to its unique application in dynamic and adaptive defect detection, particularly for real-time quality control and evolving industrial environments.
Additionally, to meet industrial demands for real-time performance and high detection efficiency, model compression techniques are frequently applied during deployment to enhance operational efficiency. While ML-based defect detection has significantly improved detection accuracy and efficiency, challenges remain, including the complexity of defect types, limited generalization of detection models, and the high cost of model training and deployment. By categorizing methods based on their application-specific relevance and unique contributions, this framework aims to provide a clear, practical, and comprehensive perspective on material defect detection technologies.
@@ 3.1. Unsupervised Learning Techniques in Material Defect Detection
This section elaborates on the UL techniques employed in MDD. Fig. 4 outlines its general workflow comprising four main steps: data collection, data preprocessing, feature extraction, and anomaly detection. Initially, a sensor captures data, forming the original dataset. This dataset is preprocessed to remove outliers, normalize values, and fill in missing data, producing a processed dataset suitable for subsequent analysis. Key features are extracted using methods such as Principal Component Analysis (PCA), histograms, and Deep Belief Networks (DBNs), which are categorized as feature-based techniques. Finally, anomaly detection is performed using techniques like clustering, DL models (e.g., VAEs and DBNs), and classification networks to identify and classify material defects.
3.1.1. Feature-Based Techniques
Feature-based techniques aim to extract meaningful and compact representations of raw data, emphasizing key characteristics that are critical for identifying defects. These methods transform complex, high-dimensional data into simplified forms, enabling subsequent detection algorithms, such as clustering or classification, to perform more efficiently and accurately. By highlighting essential features, feature-based techniques play a crucial role in enhancing the precision and robustness of defect detection workflows.
Histogram-Based
Histogram-based techniques are widely applied in MDD (kumar2021histogram, ). These methods typically analyze texture features such as colour histograms and singular value decomposition (SVD) to identify defects effectively (song2015wood, ). For instance, Weighted Object Variance (WOV) enhances Otsu’s method, improving segmentation accuracy by handling varying defect sizes and histogram distributions, resulting in higher detection rates and fewer false alarms (yuan2015improved, ). Similarly, Erazo-Aux et al. (erazo2019histograms, ) introduce a method that combines histograms of oriented gradients with pulsed thermography for defect detection in composite materials, achieving robust results with an AUC score of 0.949. Additionally, Mahmoud et al. (halfawy2014automated, ) employ a pattern recognition approach integrating HOG features and SVM classifiers for automating sewer defect detection from CCTV inspection images, demonstrating reliable segmentation and classification of defect areas.
Principal Component Analysis (PCA)-Based